A land lord has eight 40watt electric light bulbs and two 100watt bulbs in his house. If he has all the points operating for 5hrs daily and if NEPA charges 5k per unit(kWh), calculate his bill for 30 days
1
Expert's answer
2020-06-26T14:36:36-0400

The lamps consume the following amount of energy per one day:


"E_{day} = (8\cdot 40watt + 2\cdot 100watt)\cdot 5hours = 2600 Wh = 2.6kWh"

During 30 days they will consume:


"E = 30E_{day} = 78kWh"

 If NEPA charges 5k per unit(kWh), the bill will be:


"B = 5k\cdot E = 390k"

Answer. 390k